OpenStudy (anonymous):

What is the standard form of (2-4i)(3+5i)/(3+i)?

OpenStudy (anonymous):

expand the numerator and simplify, post what u get and then i'll tell u what to do next

OpenStudy (anonymous):

how do u expand?

OpenStudy (anonymous):

expand it like how you would expand a quadratic

OpenStudy (anonymous):

like foil?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok 6-2i-20i^2/3+i

OpenStudy (anonymous):

ok so what does i^2 evaluate to?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

i^2 = \[\sqrt{-1}\times \sqrt{-1}\] = -1

OpenStudy (anonymous):

yes? so what does the numerator become now?

OpenStudy (anonymous):

idk how would that change the numerator

OpenStudy (anonymous):

numerator = 6-2i-20i^2 as shown above, i^2 = -1 so 6-2i-20i^2 =6-2i-20(-1) =6-2i+20 =26-2i do you agree?

OpenStudy (anonymous):

yes i see how you did that

OpenStudy (anonymous):

so now its 26-2i/3+i

OpenStudy (anonymous):

so in math we all hate complex denominator, so we always multiply them with their conjugates to simply them what you need to do now, is evaluate: \[\frac{ (26-2i) }{ (3+i) }\times \frac{ (3-i) }{ (3-i) }\] can you do that and post your answer? make sure u simplify both the numerator and the denominator

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

38/5-16i/5

OpenStudy (anonymous):

so would the answer be 36-16i/5?

OpenStudy (anonymous):

38*

OpenStudy (anonymous):

it should be 19/2 - 4i

OpenStudy (anonymous):

does not give that answer

OpenStudy (anonymous):

the answer i put was correct but thanks for all the help to get me that far!!!

OpenStudy (anonymous):

sorry my bad, that is right, 38/5 - 16i/5

OpenStudy (anonymous):

i rechecked my working out